1. A particle with a charge of 4μC is moving at 3 X 106 m/s perpendicularly through a magnetic field with a strength of 0.10T. What is the magnitude of the force on the particle?


solution please. thanks!


1
Expert's answer
2022-05-19T05:43:07-0400
"F=qvB\sin\theta\\
=4*10^{-6}*3*10^6*0.10*\sin90^\circ=1.2\:\rm N"